OpenStudy (sawq):

Will fan and medal best help! Which of the following ordered pairs represents the solution to the system given below? 2x + y = 20 4x − 2y = 40 (10, 0) (0, 10) (4, 2) (10, −10) I know what to do I'm just not getting any of the answers given.. help?

OpenStudy (dan815):

okay well u see, you can always plug in a point to see if its working

OpenStudy (dan815):

the first point simply works (10,0) means x=10 and y=0

OpenStudy (dan815):

2x + y = 20 4x − 2y = 40 ------------- 2*10+0=20 4*10 - 2*0 = 40 ---------------- so right away (10,0) works

OpenStudy (dan815):

To actually solve this system of equations, u have to learn substitution or elimination methods, please look that up

OpenStudy (sawq):

Yeah I just learned that but its not working out...

OpenStudy (sawq):

But thanks for your help

OpenStudy (dan815):

sure

OpenStudy (dan815):

I can check your work, if u post it here, good luck